# Towards Linear Spanners in All Temporal Cliques

### v1

@@ 7 Product Graphs
Since $f(n)\in\operatorname{\mathcal{O}}(n)$, we have $s_{n}\in\Theta(n)$. By Lemma 7.8, we have that any pivot set is in $\operatorname{\mathcal{O}}(k_{n})=\operatorname{\mathcal{O}}(f(n))$, and by Lemma 7.10 any set of not reverted edges has size at least
@@
[…]
@@ 8 Conclusion
We have made a significant step forwards towards answering whether temporal cliques admit linear-size temporal spanners, see Figure 7 for an illustration. The obvious problem that stems from this paper, is whether classes of graphs exist that are not solved by our techniques. We conjecture that this is the case but the classes of graphs remaining should be quite artificial. Another possible avenue is to generalize some of our techniques to include a wider variety of graphs. For example, we could consider investigating the temporal ordering of edges and try to generalize the concept of reverted edges. Also, product graphs and their solution suggest a way in which temporal bi-cliques may be composed and decomposed. Future research could generalize this idea and more closely analyze the cases in which a temporal spanner can be obtained by combining temporal spanners on subinstances. Towards the opposite direction, it is now easier to find a $\Omega(n\log n)$ counterexample, if it exists at all, since we have provided many properties that cannot exist in a counterexample.

### v2

@@ 7 Product Graphs
Since $f(n)\in\operatorname{\mathcal{O}}(n)$, we have $s_{n}\in\Theta(n)$. By Lemma 7.8, we have that any pivot-set is in $\operatorname{\mathcal{O}}(k_{n})=\operatorname{\mathcal{O}}(f(n))$, and by Lemma 7.10 any set of not reverted edges has size at least
@@
Case 2: $a_{G}^{\prime}\neq\pi^{+}_{{S_{G}}}({b_{G}})$. We append $\mathopen{}\mathclose{{}\left\{{(a_{G}^{\prime},a_{H})},{(b_{G},b_{H}^{*})}}\right\}$, which is in $S$ because $\mathopen{}\mathclose{{}\left\{{a_{G}^{\prime}},{b_{G}}}\right\}$ is the last edge of $P_{G}$. Further, we can append $\mathopen{}\mathclose{{}\left\{{(\pi^{+}_{{S_{G}}}({b_{G}}),a_{H}^{*})},{(b_{G},b_{H}^{*}}}\right\}$ to $P$, since $\mathopen{}\mathclose{{}\left\{{\pi^{+}_{{S_{G}}}({b_{G}})},{b_{G}}}\right\}\in S_{G}$. Again, as $S_{H}$ is a bi-spanner for $H$, there exists a temporal path $P^{\prime}_{H}$ from $a_{H}^{*}$ to $b_{H}$ in $S_{H}$. In the same fashion as in Case 1, for every edge $\mathopen{}\mathclose{{}\left\{{a_{H}^{P}},{b_{H}^{P}}}\right\}$ in $P^{\prime}_{H}$, we append $\mathopen{}\mathclose{{}\left\{{(\pi^{+}_{{S_{G}}}({b_{G}}),a_{H}^{P})},{(b_{G},b_{H}^{P})}}\right\}$ to $P$. See Figure 6(b) for a visualization of the constructed path $P$. As $\mathopen{}\mathclose{{}\left\{{a_{H}^{*}},{b_{H}^{*}}}\right\}$ has the minimal label in $S_{H}$ and $a_{G}^{\prime}\neq\pi^{+}_{{S_{G}}}({b_{G}})$, our construction yields a temporal path which is clearly included in $S$.
@@ 8 Conclusion
We present a significant step forwards towards answering whether temporal cliques admit linear-size temporal spanners, see Figure 7 for an illustration. The obvious problem that stems from this paper, is whether classes of graphs exist that are not solved by our techniques. We conjecture that this is the case but the classes of graphs remaining should be quite artificial. Another possible avenue is to generalize some of our techniques to include a wider variety of graphs. For example, we could consider investigating the temporal ordering of edges and try to generalize the concept of reverted edges. Also, product graphs and their solution suggest a way in which temporal bi-cliques may be composed and decomposed. Future research could generalize this idea and more closely analyze the cases in which a temporal spanner can be obtained by combining temporal spanners on subinstances. Towards the opposite direction, it is now easier to find a $\Omega(n\log n)$ counterexample, if it exists at all, since we have provided many properties that any counterexample must avoid.